Ralph and Adam earn $600 cleaning a huge house. Adam was the one who found the job,so he took 60%of the earnings. How much Maloney did Ralph take

Answer by akch2002(12)

Total earnings = $600
Adam's share = 60%
Hence, Ralph's share = 100-60 = 40%
Hence, Ralph's earning = 40% of 600
= (40/100)*600
= $240